32. An air-conditioning system is used to maintain a house at 65oF when the temperature outside is 85oF. The house is gaining heat through the walls and windows at a rate of 1250 Btu/min, and the heat generation rate within the house from people, lights, and appliances amounts to 350 Btu/min. Determine the minimum power input required for this air-conditioning system. A


1
Expert's answer
2022-03-16T01:26:05-0400

Solution;

"T_L=65\u00b0F"

"T_H=85\u00b0F"

"Q=1250+350=1600Btu\/min"

min

"COP=\frac{1}{(\frac{T_H}{T_L})-1}"

"COP=\frac{1}{(\frac{85+460}{65+460})-1}=26.25"

"P_{in}=\frac{Q}{COP}=\frac{1600}{26.25}=60.95Btu/min"